if the concentration of A is doubled and the concentration of B is halved, what happens to the reaction rate
1
Expert's answer
2019-11-05T07:39:47-0500

If we have reaction: A + B = AB

The reaction rate from the Law of mass action is: V(1) = k*[A]*[B]

if the concentration of A is doubled(2*[A]) and the concentration of B is halved(1/2*[B]), the reaction rate become: V(2) = k*2*[A]*1/2*[B] = k*[A]*[B].

So reaction rate doesn't change: V(1)=V(2).